# Inflated hot Jupiters: inferring average atmospheric velocity via Ohmic models coupled with internal dynamo evolution

### v1

@@ 5.2 Intermittent dynamo and its feedback on the atmospheric induction
For highly irradiated and highly heated models, the convection region can be reduced in size, or broken into two or more thin convective shell. This can happen when most of the heat is deposited in the outermost layers of the original convective region, so that the normal temperature gradient is reduced, or even inverted. In fact, this has been noticed in past works which deposited heat in outer layers only (Komacek & Youdin, 2017; Komacek et al., 2020; Elias-López et al., 2025a), or used Ohmic heating profiles strongly decaying towards the interior (Batygin & Stevenson, 2010; Batygin et al., 2011; Wu & Lithwick, 2013; Knierim et al., 2022), similar to what is derived here. In those earlier studies, the consequence of having alternate convective and radiative interior layers was to delay cooling. However, they assumed no feedback on the deposited heat, i.e., they kept a constant-in-time Ohmic efficiency.
@@
In our model, a decrease of the convective region implies a lower average heat flux, which implies a lower dynamo magnetic field, via eq. (18). As a consequence, the deposited heat, $\propto B_{\rm atm}^{2}$, decreases. Therefore, in time, the convection across a very thick shell (most of the planet) is allowed again, which rises $B_{\rm atm}$. This mechanism implies that the feedback between background field $B_{\rm atm}(L)$ and the Ohmic dissipation rate $Q_{j}(B_{\rm atm})$, generates an unstable mechanism that leads to cyclic rises and falls of the heat flux, and therefore of the Ohmic dissipation. This is visible in several models of Fig. 6.
@@
We illustrate this effect in Fig. 7, which considers two representative models with the same $M=1\leavevmode\nobreak\ M_{j}$, $T_{\rm eq}=1750$ K and $v_{\rm avg}=793$ m/s. One is the standard model considered in this paper, with an evolving $B_{\rm atm}$ coupled to the Ohmic heating (coupled model, grey lines); in the other, we manually fix throughout the evolution $B_{\rm atm}=5$ G (a value which is in the range of the evolving $B_{\rm atm}$ case at Gyr; the results would be exactly the same for any given product $B_{\rm atm}v_{\rm avg}$). In the left panel, we show the values of the pressure delimiting the convective regions. The middle panel shows the surface (dots), Ohmic (dashes), and internal luminosity (solid). The latter is evaluated at $p=p_{\rm dyn}$, considering the dynamo region and not the Ohmic contributions in the less-conductive envelope where no dynamo is present, and is a tracer of the convective heat flux $q(r)$ entering in the scaling laws for $B_{\rm dyn}$ (and, consequently, $B_{\rm atm}$). The right panel shows the difference in the planetary and dynamo radii evolution.
@@
We can relate the results illustrated here to the cooling regimes presented by Komacek & Youdin (2017) and Ginzburg & Sari (2016), which correspond to different relative weights of irradiation power, internal luminosity, and total heat rate ($L_{\rm j}$ in our case). Heated HJs generally fall in the case for which $L_{\rm int}\ll L_{\rm j}\ll L_{\rm irr}$, labelled as regime 2 in Komacek & Youdin (2017) and stages 3 or 4 in Ginzburg & Sari (2016). Within this regime, there are qualitative differences, depending on the location and intensity of the heat source. In our models, the heat is always spread continuously over the radiative and convective regions, see Fig. 3. If the heating is confined to radiative layers, with negligible amount in the convective region, there is a perpetual cooling (regime 2d of Komacek & Youdin 2017), with a progressive deepening of the radiative-convective boundary (RCB), since the internal luminosity decreases in time while the irradiation power stays constant (or increases if the stellar luminosity evolves, see below). However, in many of the models shown here, the relevant scenario is the appearance of a shallow convective region, and a radiative layer just below the region where the bulk of heat is deposited. Such transition, already identified as appropriate for heated HJs by Komacek & Youdin (2017), is clearly seen in the fixed $B_{\rm atm}$ case (red lines) of Fig. 7, left panel, described as follows. Before $\sim 0.4$ Gyr, there is only one convective region extending up to tens of bars (regime 2b of Komacek & Youdin 2017). As the internal luminosity decreases, it becomes comparable or lower than the amount of deposited heat inside the convective region, which is a fraction of the $L_{\rm j}$ indicated in the middle panel: therefore, a radiative region opens around $p\sim 5000$ bar, and slowly expands, splitting the convection in two separated layers below and above it. Physically, this happens because the additional heat, which is concentrated in the outermost regions (Fig. 3), lowers the local temperature gradient to sub-adiabatic values in a region that is typically located just beneath the main portion of the heated area.
@@
For the coupled model (grey lines in Fig. 7), however, the scenario is more complicated due to the sudden changes of the heat power. We stress that an exact comparison with Komacek & Youdin (2017) is not possible in this case for two reasons. The first, valid also for the fixed $B_{\rm atm}$ case, is that they considered a localized heat source at a given pressure, entirely in the convective region, while ours extends from the surface down to the convective region, more similar to the power laws of Ginzburg & Sari (2016), who however did not consider the relevant regime 2b of Komacek & Youdin (2017). The second is that their heat source is constant in time, while ours vary due to the background field coupling (except in the fixed $B_{\rm atm}$ case shown above). Qualitatively, we can affirm that the relevant regimes are the same as the ones described above (2b and 2c), but with multiple transitions between them. This cyclic behaviour works as follows: when the heat flux is reduced due to the broken convection, $B_{\rm dyn}$, $B_{\rm atm}$ and the Ohmic power $L_{j}$ also decreases, although the instantaneous variations are averaged out in time, as described in Sect. 4.4. Since the heating is less, the full convection can be restored. This trend is repeated many times, although the time average of the Ohmic source affects the duration of the oscillations and we don’t claim of having properly captured the mechanism. We cannot exclude that the numerical difficulties in converging to a solution contribute to the short-term fluctuations, visible as noise in Ohmic efficiency and radius evolution. However, the qualitative difference between the two models is clear.
@@
Another important difference between the fixed $B_{\rm atm}$ and coupled models is seen when comparing the radius and Ohmic efficiency evolution. At contrast with the coupled models, the Ohmic efficiency is practically constant in the fixed $B_{\rm atm}$ case, the only change can be due to $\sigma_{\rm atm}$, see eq. (10). Consequently, the radius varies much more slowly and maintains larger values, similarly to the models with constant efficiency present in the literature (e.g., Batygin et al. 2011; Komacek & Youdin 2017; Thorngren & Fortney 2018; Knierim et al. 2022). This reinforces the peculiarity and complexity of the Ohmic models here presented, considering the evolution of $B_{\rm atm}$ and its coupling to the cooling.
@@ 5.3 Effects of evolving the stellar luminosity
Note that the luminosity evolution significantly depends on the host star type. Hence its inclusion effectively breaks the degeneracy between host star type and orbital period, which has been implicit in all the results so far (different combinations of them provide the same irradiation or $T_{\rm eq}$). Therefore, since the focus of the paper is rather on the effect of coupling the dynamo field to the induced currents, here we only consider a few representative cases, for $1\leavevmode\nobreak\ M_{j}$ HJs orbiting a Sun-like star. In Fig. 8, we show the evolution of radius, Ohmic efficiency and estimated surface dipolar magnetic fields for cases of evolving stellar luminosities (dashed lines) having $T_{\rm eq}=1500$ K and $2000$ K at $t=5$ Gyr, and we compare them with the cases with the same fixed $T_{\rm eq}$ (solid lines) shown in the previous sections. We show both the cases without and with heating, for a representative value of $v_{\rm avg}$. The evolution of the radius indicates that, indeed, re-inflation at $t\gtrsim 5$ Gyr is possible. In fact, the overall trends, such as the decline in Ohmic efficiencies, magnetic field strengths, and the on/off mechanism of the internal convective heat determining the dynamo, are not qualitatively altered by the inclusion of the evolving luminosity during the first few Gyr. However, during the later stages of the main sequence, the increasing luminosity plays a significant role: it can halt planetary contraction even in the absence of internal heating, purely through an enhanced blanketing effect. When heating is present, this impact becomes dramatically more pronounced. Moreover, we notice how the heated cases with $T_{\rm eq}(t=5\,{\rm Gyr})=1500$ K or 2000 K behave differently: in the former, the Ohmic efficiency increases after $\sim 4$ Gyr, while in the latter it drops faster than in the corresponding fixed $T_{\rm eq}$ case. This is due, again, to the powerful effect of suppressing convection and dynamo in large parts of the interior in case of highly irradiated and highly heated HJs, as evident also from the values of $B_{\rm atm}$ derived from the convective flux (bottom panel of Fig. 8).
@@
A thorough investigation of this effect warrants a systematic exploration of the relevant parameters, with the stellar mass (i.e., spectral type) as an additional parameter. Moreover, for the sake of self-consistency, one should also account for the variation of $v_{\rm avg}$, since it is not realistic to allow evolution of the irradiation and of the internal magnetic field while keeping constant the winds. Addressing this properly is beyond the scope of this work and should be pursued in future studies. This section is intended merely to offer a preliminary glimpse into the potentially complex, non-trivial effects of an evolving stellar luminosity.
@@
Figure 8: Comparison of the planetary radii (top), Ohmic efficiency (middle), and background magnetic field at the surface $B_{\rm atm}$ (bottom), between models with fixed (solid lines) and varying (dashed) $T_{\rm eq}$. We consider models with $T_{\rm eq}=1500$ K (black and blue lines with $v_{\rm avg}=0$ and 1000 m/s, respectively) and 2000 K (grey and red lines with $v_{\rm avg}=0$ and 178 m/s, respectively), where these values of $T_{\rm eq}$ are at $t=5$ Gyr in the cases with evolving stellar luminosity.
@@ 5.4 Implications for the internal dynamo region
For comparison, and to get an insight into the possible interplay between atmospheric and deep-seated dynamos, one can compare the amount of atmospherically-induced currents in deep layers with the ones independently generated by the active dynamo. In dynamo simulations, it is not possible to directly predict the absolute value of the magnetic field and currents, since they depend on dynamo parameters like the Rayleigh, Ekman and Prandtl numbers, which are computationally constrained to be many orders of magnitude away from the realistic ones (due to the impossibility of including all relevant scales). However, we can use the information on the typical topology from dynamo simulations, by looking at the ratio between currents and magnetic fields, and re-scale them by using the magnetospheric values measured in the solar planets. In particular, here we make use of the observations of the Jovian field, the one reconstructed with the highest fidelity among gas giants. We have considered the Juno measurements of the Jovian magnetic field above the surface, which is represented in terms of multipolar decomposition of the poloidal and toroidal components, as given by the best-fit model to the data (Connerney et al., 2018, 2022). With the set of these best-fitting weights, one can describe the potential magnetic field in the entire volume where there are no currents, down to the radius $R_{\rm dyn}$, below which the dynamo starts to be active and the field is no longer potential. From this, one can then calculate the field line curvature vector, defined by $\boldsymbol{\kappa}=(\boldsymbol{b}\cdot\boldsymbol{\nabla})\boldsymbol{b}$, where $\boldsymbol{b}\equiv\boldsymbol{B}/B$. We consider the radial profile, taking spatial averages over a sphere at a given radius, $\bar{\kappa}$. In App. B.1, we show that, for Jupiter, the inferred values are $\bar{\kappa}\sim(0.07-0.3\leavevmode\nobreak\ R_{j})^{-1}$, depending on the magnetospheric model used and the depth considered. We adopt this as a lower limit for $\bar{\kappa}$, since higher values are expected in deeper regions, as a consequence of the turbulent environment in which the magnetic field is continuously regenerated (see bottom panel of Fig. 9).
@@
Figure 9: Typical radial profiles of |×$|\boldsymbol{\nabla}\times\boldsymbol{B}|/(\bar{\kappa}B)$ (top) and $1/\bar{\kappa}$ (bottom), averaged over the spherical surface at each radius, from 3D dynamo simulations with the code MagIC. In order to explore the sensitivity to the free parameters. All models have the same boundary conditions (i.e. stress-free for the fluid and insulating for the magnetic fields outside, stress-free and perfectly conducting inside) except for the cases marked in the legend. The Ekman number is $10^{-5}$, and both Prandtl numbers are set to 1 (except for those also marked in the legend, which have $N_{\rho}\approx 3.0$ and ${\rm Ra}=5\times 10^{8}$. Note that $1/\bar{\kappa}$ becomes much smaller close to the external boundary, in order to match with the potential configuration constraint. See App. B.2 for more details.
@@
While we can infer the minimum value of $\bar{\kappa}$ from Juno measurements, we can relate it to the values of currents $J$ by looking at simulations. To do so, we performed MHD dynamo simulations that allowed us to estimate the values of the currents, and to validate the use of the magnetospheric value $\bar{\kappa}$ as a proxy for the internal $J/B$ ratio. We have used the code MagIC (Jones et al., 2011; Gastine & Wicht, 2012) to study the amplification of the magnetic field in HJ-like planets. We used a variety of dynamo parameters and boundary conditions, and a representative structure profile for conductivity, density and other thermodynamic quantities, coming from our irradiated HJ MESA models, following the methodology of Elias-López et al. (2025b) (see App. B.2 for details). These simulations provide us with a self-consistent evolution of the magnetic field as amplified in the convective, turbulent interior of a HJ, and, therefore, allow a robust estimation of the typical length scale $\kappa$. Note that a separate upcoming work will focus particularly on the 3D dynamo solutions for HJs. Here, we only focus on a specific aspect, which helps us to qualitatively compare the intensity of atmospherically induced and dynamo currents.
@@
We show the results in Fig. 9, where only some representative models producing a dynamo are shown. Despite some small differences among the various simulations, we obtain radial profiles with the same trend and order of magnitude, both for |×$|\boldsymbol{\nabla}\times\boldsymbol{B}|/(\bar{\kappa}B)\sim 1$-$3$ and $1/\bar{\kappa}\sim(0.01$-$0.1)\leavevmode\nobreak\ R_{j}$. For the former (top panel), it confirms that the ratio between the magnetic field intensity and the mean radius of curvature, $B\bar{\kappa}$, is a very good tracer of the current density, $J$. Regarding $\bar{\kappa}$ itself (bottom panel), we obtain values similar to the ones inferred from Juno data (App. B.1). More realistic outer boundary conditions for HJs might potentially alter the curvature $\bar{\kappa}$ near the external layers, but as seen in Fig. 9, the various magnetic boundary conditions tested have no significant impact on the values within the bulk of the simulation.
@@
Therefore, taking typical values $B_{\rm dyn}\sim 10$ G and $\kappa\sim 10^{-6}$ m-1, one obtains $J\sim B\kappa/\mu_{0}\sim 10^{-3}$ A/m2. This range is the same as (or slightly lower than) the atmospherically-induced currents at $r\sim R_{\rm dyn}$ (see top panel of Fig. 3). Therefore, they could be a significant factor to be taken into account to set the boundary conditions in dynamo simulations, which are instead almost always set to be a current-free solution. Since the geometry of the atmospherically induced currents depends on the wind and background field configurations themselves, this opens to the remarkable possibility of atmospheric processes affecting the deep dynamo solution.

### v2

@@ 5.2 Intermittent dynamo and its feedback on the atmospheric induction
For highly irradiated and highly heated models, the convection region can be reduced in size, or broken into two or more thin convective shell. This can happen when most of the heat is deposited in the outermost layers of the original convective region, so that the normal temperature gradient is reduced, or even inverted. In fact, this has been noticed in past works which deposited heat in outer layers only (Komacek & Youdin, 2017; Komacek et al., 2020; Elias-López et al., 2025a), or used Ohmic heating profiles strongly decaying towards the interior (Batygin & Stevenson, 2010; Batygin et al., 2011; Wu & Lithwick, 2013; Knierim et al., 2022), similar to what is derived here. In those earlier studies, the consequence of having alternate convective and radiative interior layers was to delay cooling. However, they assumed no feedback on the deposited heat, i.e. they kept a constant-in-time Ohmic efficiency.
@@
In our model, a decrease of the convective region implies a lower average heat flux, which implies a lower dynamo magnetic field, via eq. (18). As a consequence, the deposited heat, $\propto B_{\rm atm}^{2}$, decreases. Therefore, in time, the convection across a very thick shell (most of the planet) is allowed again, which rises $B_{\rm atm}$. This mechanism implies that the feedback between background field $B_{\rm atm}(L)$ and the Ohmic dissipation rate $Q_{\rm j}(B_{\rm atm})$, generates an unstable mechanism that leads to cyclic rises and falls of the heat flux and therefore of the Ohmic dissipation. This is visible in several models of Fig. 6.
@@
We illustrate this effect in Fig. 7, which considers two representative models with the same $M=1\leavevmode\nobreak\ M_{\rm j}$, $T_{\rm eq}=1750$ K and $v_{\rm avg}=793$ m/s. One is the standard model considered in this paper, with an evolving $B_{\rm atm}$ coupled to the Ohmic heating (coupled model, grey lines); in the other, we manually fix throughout the evolution $B_{\rm atm}=5$ G (a value which is in the range of the evolving $B_{\rm atm}$ case at gigayears; the results would be exactly the same for any given product $B_{\rm atm}v_{\rm avg}$). In the left panel, we show the values of the pressure delimiting the convective regions. The middle panel shows the surface (dots), Ohmic (dashes), and internal luminosity (solid). The latter is evaluated at $p=p_{\rm dyn}$, considering the dynamo region and not the Ohmic contributions in the less-conductive envelope where no dynamo is present, and is a tracer of the convective heat flux $q(r)$ entering in the scaling laws for $B_{\rm dyn}$ (and, consequently, $B_{\rm atm}$). The right panel shows the difference in the planetary and dynamo radii evolution.
@@
We can relate the results illustrated here to the cooling regimes presented by Komacek & Youdin (2017) and Ginzburg & Sari (2016), which correspond to different relative weights of irradiation power, internal luminosity, and total heat rate ($L_{\rm j}$ in our case). Heated HJs generally fall in the case for which $L_{\rm int}\ll L_{\rm j}\ll L_{\rm irr}$, labelled as regime 2 in Komacek & Youdin (2017) and stages 3 or 4 in Ginzburg & Sari (2016). Within this regime, there are qualitative differences, depending on the location and intensity of the heat source. In our models, the heat is always spread continuously over the radiative and convective regions, see Fig. 3. If the heating is confined to radiative layers, with negligible amount in the convective region, there is a perpetual cooling (regime 2d of Komacek & Youdin 2017), with a progressive deepening of the radiative-convective boundary (RCB), since the internal luminosity decreases in time while the irradiation power stays constant (or increases if the stellar luminosity evolves, see below). However, in many of the models shown here, the relevant scenario is the appearance of a shallow convective region, and a radiative layer just below the region where the bulk of heat is deposited. Such a transition, already identified as appropriate for heated HJs by Komacek & Youdin (2017), is clearly seen in the fixed $B_{\rm atm}$ case (red lines) of Fig. 7, left panel, described as follows. Before $\sim 0.4$ Gyr, there is only one convective region extending up to tens of bars (regime 2b of Komacek & Youdin 2017). As the internal luminosity decreases, it becomes comparable or lower than the amount of deposited heat inside the convective region, which is a fraction of the $L_{\rm j}$ indicated in the middle panel: therefore, a radiative region opens around $p\sim 5000$ bar, and slowly expands, splitting the convection in two separated layers below and above it. Physically, this happens because the additional heat, which is concentrated in the outermost regions (Fig. 3), lowers the local temperature gradient to sub-adiabatic values in a region that is typically located just beneath the main portion of the heated area.
@@
For the coupled model (grey lines in Fig. 7), however, the scenario is more complicated due to the sudden changes of the heat power. We stress that an exact comparison with Komacek & Youdin (2017) is not possible in this case for two reasons. The first reason, valid also for the fixed $B_{\rm atm}$ case, is that they considered a localised heat source at a given pressure entirely in the convective region, while our heat deposition extends from the surface down to the convective region and is more similar to the power laws of Ginzburg & Sari (2016), who, however, did not consider the relevant regime 2b of Komacek & Youdin (2017). The second reason is that their heat source is constant in time, while ours varies due to the background field coupling (except in the fixed $B_{\rm atm}$ case shown above). Qualitatively, we can affirm that the relevant regimes are the same as the ones described above (2b and 2c) but with multiple transitions between them. This cyclic behaviour works as follows: When the heat flux is reduced due to the broken convection, $B_{\rm dyn}$, $B_{\rm atm}$ and the Ohmic power, $L_{\rm j}$, also decreases, although the instantaneous variations are averaged out in time, as described in Sect. 4.4. Since the heating is lower, the full convection can be restored. This trend is repeated many times, although the time average of the Ohmic source affects the duration of the oscillations, and we do not claim to have properly captured the mechanism. We cannot exclude that the numerical difficulties in converging to a solution contribute to the short-term fluctuations, visible as noise in Ohmic efficiency and radius evolution. However, the qualitative difference between the two models is clear.
@@
Another important difference between the fixed $B_{\rm atm}$ and coupled models is seen when comparing the radius and Ohmic efficiency evolution. At contrast with the coupled models, the Ohmic efficiency is practically constant in the fixed $B_{\rm atm}$ case, the only change can be due to $\sigma_{\rm atm}$, see eq. (10). Consequently, the radius varies much more slowly and maintains larger values, similarly to the models with constant efficiency present in the literature (e.g. Batygin et al. 2011; Komacek & Youdin 2017; Thorngren & Fortney 2018; Knierim et al. 2022). This reinforces the peculiarity and complexity of the Ohmic models here presented, considering the evolution of $B_{\rm atm}$ and its coupling to the cooling.
@@ 5.3 Effects of evolving the stellar luminosity
We note that the luminosity evolution significantly depends on the host star type. Hence its inclusion effectively breaks the degeneracy between host star type and orbital period, which has been implicit in all the results so far (different combinations of them provide the same irradiation or $T_{\rm eq}$). Therefore, since the focus of the paper is rather on the effect of coupling the dynamo field to the induced currents, here we only consider a few representative cases, for $1\leavevmode\nobreak\ M_{\rm j}$ HJs orbiting a Sun-like star. In Fig. 8, we show the evolution of radius, Ohmic efficiency and estimated surface dipolar magnetic fields for cases of evolving stellar luminosities (dashed lines) having $T_{\rm eq}=1500$ K and $2000$ K at $t=5$ Gyr, and we compare them with the cases with the same fixed $T_{\rm eq}$ (solid lines) shown in the previous sections. We show both the cases without and with heating, for a representative value of $v_{\rm avg}$. The evolution of the radius indicates that, indeed, re-inflation at $t\gtrsim 5$ Gyr is possible. In fact, the overall trends, such as the decline in Ohmic efficiencies, magnetic field strengths, and the on/off mechanism of the internal convective heat determining the dynamo, are not qualitatively altered by the inclusion of the evolving luminosity during the first few gigayears. However, during the later stages of the main sequence, the increasing luminosity plays a significant role: it can halt planetary contraction even in the absence of internal heating, purely through an enhanced blanketing effect. When heating is present, this impact becomes dramatically more pronounced. Moreover, we noticed how the heated cases with $T_{\rm eq}(t=5\,{\rm Gyr})=1500$ K or 2000 K behave differently: In the former, the Ohmic efficiency increases after $\sim 4$ Gyr, while in the latter it drops faster than in the corresponding fixed $T_{\rm eq}$ case. This is due, again, to the powerful effect of suppressing convection and dynamo in large parts of the interior in case of highly irradiated and highly heated HJs, as evident also from the values of $B_{\rm atm}$ derived from the convective flux (bottom panel of Fig. 8).
@@
A thorough investigation of this effect warrants a systematic exploration of the relevant parameters, with the stellar mass (i.e. spectral type) as an additional parameter. Moreover, for the sake of self-consistency, one should also account for the variation of $v_{\rm avg}$, since it is not realistic to allow evolution of the irradiation and of the internal magnetic field while keeping constant the winds. Addressing this properly is beyond the scope of this work and should be pursued in future studies. This section is intended merely to offer a preliminary glimpse into the potentially complex, non-trivial effects of an evolving stellar luminosity.
@@
Figure 8: Comparison of the planetary radii (top), Ohmic efficiency (middle), and background magnetic field at the surface $B_{\rm atm}$ (bottom) between models with a fixed (solid lines) and varying (dashed) $T_{\rm eq}$. We consider models with $T_{\rm eq}=1500$ K (black and blue lines with $v_{\rm avg}=0$ and 1000 m/s, respectively) and 2000 K (grey and red lines with $v_{\rm avg}=0$ and 178 m/s, respectively), where these values of $T_{\rm eq}$ are at $t=5$ Gyr in the cases with evolving stellar luminosity.
@@ 5.4 Implications for the internal dynamo region
For comparison, and to get an insight into the possible interplay between atmospheric and deep-seated dynamos, one can compare the amount of atmospherically induced currents in deep layers with the ones independently generated by the active dynamo. In dynamo simulations, it is not possible to directly predict the absolute value of the magnetic field and currents, since they depend on dynamo parameters like the Rayleigh, Ekman and Prandtl numbers, which are computationally constrained to be many orders of magnitude away from the realistic ones (due to the impossibility of including all relevant scales). However, we can use the information on the typical topology from dynamo simulations, by looking at the ratio between currents and magnetic fields, and re-scale them by using the magnetospheric values measured in the solar planets. In particular, here we make use of the observations of the Jovian field, the one reconstructed with the highest fidelity among gas giants. We have considered the Juno measurements of the Jovian magnetic field above the surface, which is represented in terms of multipolar decomposition of the poloidal and toroidal components, as given by the best-fit model to the data (Connerney et al., 2018, 2022). With the set of these best-fitting weights, one can describe the potential magnetic field in the entire volume where there are no currents, down to the radius $R_{\rm dyn}$, below which the dynamo starts to be active and the field is no longer potential. From this, one can then calculate the field line curvature vector, defined by $\boldsymbol{\kappa}=(\boldsymbol{b}\cdot\boldsymbol{\nabla})\boldsymbol{b}$, where $\boldsymbol{b}\equiv\boldsymbol{B}/B$. We consider the radial profile, taking spatial averages over a sphere at a given radius, $\bar{\kappa}$. In App. B.1, we show that, for Jupiter, the inferred values are $\bar{\kappa}\sim(0.07-0.3\leavevmode\nobreak\ R_{\rm j})^{-1}$, depending on the magnetospheric model used and the depth considered. We adopt this as a lower limit for $\bar{\kappa}$, since higher values are expected in deeper regions, as a consequence of the turbulent environment in which the magnetic field is continuously regenerated (see bottom panel of Fig. 9).
@@
Figure 9: Typical radial profiles of |×$|\boldsymbol{\nabla}\times\boldsymbol{B}|/(\bar{\kappa}B)$ (top) and $1/\bar{\kappa}$ (bottom), averaged over the spherical surface at each radius, from 3D dynamo simulations with the code MagIC. In order to explore the sensitivity to the free parameters. All models have the same boundary conditions (i.e. stress-free for the fluid and insulating for the magnetic fields outside, stress-free and perfectly conducting inside) except for the cases marked in the legend. The Ekman number is $10^{-5}$, and both Prandtl numbers are set to 1 (except for those also marked in the legend, which have $N_{\rho}\approx 3.0$ and ${\rm Ra}=5\times 10^{8}$. We note that $1/\bar{\kappa}$ becomes much smaller close to the external boundary, in order to match with the potential configuration constraint. See App. B.2 for more details.
@@
While we can infer the minimum value of $\bar{\kappa}$ from Juno measurements, we can relate it to the values of currents $J$ by looking at simulations. To do so, we performed MHD dynamo simulations that allowed us to estimate the values of the currents, and to validate the use of the magnetospheric value $\bar{\kappa}$ as a proxy for the internal $J/B$ ratio. We have used the code MagIC (Jones et al., 2011; Gastine & Wicht, 2012) to study the amplification of the magnetic field in HJ-like planets. We used a variety of dynamo parameters and boundary conditions, and a representative structure profile for conductivity, density and other thermodynamic quantities, coming from our irradiated HJ MESA models, following the methodology of Elias-López et al. (2025b) (see App. B.2 for details). These simulations provide us with a self-consistent evolution of the magnetic field as amplified in the convective, turbulent interior of a HJ and therefore allow for a robust estimation of the typical length scale $\kappa$. We note that a separate upcoming work will focus particularly on the 3D dynamo solutions for HJs. Here, we only focus on a specific aspect, which helps us to qualitatively compare the intensity of atmospherically induced and dynamo currents.
@@
We show the results in Fig. 9, where only some representative models producing a dynamo are shown. Despite some small differences among the various simulations, we obtain radial profiles with the same trend and order of magnitude, both for |×$|\boldsymbol{\nabla}\times\boldsymbol{B}|/(\bar{\kappa}B)\sim 1$-$3$ and $1/\bar{\kappa}\sim(0.01$-$0.1)\leavevmode\nobreak\ R_{\rm j}$. For the former (top panel), it confirms that the ratio between the magnetic field intensity and the mean radius of curvature, $B\bar{\kappa}$, is a very good tracer of the current density, $J$. Regarding $\bar{\kappa}$ itself (bottom panel), we obtain values similar to the ones inferred from Juno data (App. B.1). More realistic outer boundary conditions for HJs might potentially alter the curvature $\bar{\kappa}$ near the external layers, but as seen in Fig. 9, the various magnetic boundary conditions tested have no significant impact on the values within the bulk of the simulation.
@@
Therefore, taking typical values $B_{\rm dyn}\sim 10$ G and $\kappa\sim 10^{-6}$ m-1, one obtains $J\sim B\kappa/\mu_{0}\sim 10^{-3}$ A/m2. This range is the same as (or slightly lower than) the atmospherically induced currents at $r\sim R_{\rm dyn}$ (see top panel of Fig. 3). Therefore, they could be a significant factor to be taken into account to set the boundary conditions in dynamo simulations, which are instead almost always set to be a current-free solution. Since the geometry of the atmospherically induced currents depends on the wind and background field configurations themselves, this opens to the remarkable possibility of atmospheric processes affecting the deep dynamo solution.